You are the operations manager for an airline and you are considering a higher fare level for passengers in aisle seats. How many randomly selected air passengers must you​ survey? Assume that you want to be 95​% confident that the sample percentage is within 4.5 percentage points of the true population percentage. Complete parts​ (a) and​ (b) below. a. Assume that nothing is known about the percentage of passengers who prefer aisle seats. nequals 475 ​(Round up to the nearest​ integer.) b. Assume that a prior survey suggests that about 39​% of air passengers prefer an aisle seat.
Answer #1

SOLUTION:

Solution :

Given that,

= 0.39

1 - = 1 - 0.39 = 0.61

margin of error = E = 4.5% = 0.045

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

Sample size = n = (Z/2 / E)2 * * (1 - )

= (1.96 / 0.045)2 * 0.39 * 0.61

= 451.31

Sample size = 452 (ROUNDED)